# Transfer Matrix and Lattice Dilatation Operator for High-Quality Fixed Points in Tensor Network Renormalization Group

### v1

@@ 2.2 Tensor RG
An RG map $\mathcal{R}:A\to A^{\prime}$ preserves the partition function while simultaneously reducing the dimensions of the network by a constant factor $b>1$:
@@ 2.4 Critical fixed point via Newton method
where $P_{s}$ is the orthogonal projector on the subspace spanned by the first $s$ $\mathbb{Z}_{2}$-even (see footnote 3) eigenvectors of $\nabla R^{\circ}(A_{(0)})$. For not too large $s$, these eigenvectors are affordable to compute using iterative Krylov subspace methods.
@@ 3.1 Theory
We will need to consider tensor network partition functions with twisted boundary conditions. We denote by $Z_{h,v}(A,N\times M)$ the partition function of $N\times M$ tensor network built out of tensor $A$ such that the bottom legs $i$ are contracted with the top legs $i+h$ modulo $N$, and left legs $j$ are contracted with right legs $j+v$ modulo $M$:
@@
For $h=v=0$ we recover the usual partition function with the periodic boundary conditions.
@@
Let $\mathcal{R}$ be a non-rotating RG map with scale factor $b$, and let $h,v$ be two nonnegative integers divisible by $b$. We will say that $\mathcal{R}$ is $(h,v)$-twist friendly, if it preserves the twisted partition function up to reducing all dimensions and twists by the same factor $b$. I.e., for $N,M$ divisible by $b$, we have:
@@
The twist friendliness condition is needed for the validity of some TM spectrum properties stated below. All RG maps known to us, and in particular Gilt-TNR, are $(h,v)$-twist friendly for any $h,v$ divisible by $b$. Intuitively, twist friendliness is a consequence of “locality” of the RG map, and it appears to have been tacitly assumed in TEFR . Rather than trying to formalize the connection to locality, we prefer to state this condition explicitly.
@@ 4.1 LDO definition
Some remarks are in order here. First, note that we posit the existence of an operator $\mathcal{D}$ which acts from the networks with $2\times 2$ holes to networks with $2\times 2$ holes. It’s not obvious that the hole shape may be so preserved. Instead, it can happen that the hole shrinks or expands in size. The shrinking hole case will be discussed in Appendix C, where we will explain some aspects of the results of Lyu:2021qlw ; PhysRevE.109.034111 (see footnote 5). Here in the main text we focus on the $2\times 2$ hole preserving case since this turns out to be relevant for Gilt-TNR.

### v2

@@ 2.2 Tensor RG
An RG map $\mathcal{R}:A\to A^{\prime}$ preserves the partition function while simultaneously reducing the dimensions of the network by a constant integer factor $b>1$: [^2] We note that a single step of the TRG map on square lattice, (Levin:2006jai, , Fig. 3), would not fit this definition as it rotates the lattice by $\pi/4$ and reduces the dimensions by $\sqrt{2}$. Two steps of TRG would be OK, with $b=2$.
@@ 2.4 Critical fixed point via Newton method
where $P_{s}$ is the orthogonal projector on the subspace spanned by the first $s$ $\mathbb{Z}_{2}$-even (see footnote 4) eigenvectors of $\nabla R^{\circ}(A_{(0)})$. For not too large $s$, these eigenvectors are affordable to compute using iterative Krylov subspace methods.
@@ 3.1 Theory
Let $h,v$ be two integers. We will need to consider tensor network partition functions with twisted boundary conditions. We denote by $Z_{h,v}(A,N\times M)$ the partition function of $N\times M$ tensor network built out of tensor $A$ such that the bottom legs $i$ are contracted with the top legs $i+h$ modulo $N$, and left legs $j$ are contracted with right legs $j+v$ modulo $M$:
@@
When only one of two numbers $h,v$ is nonzero, the contraction procedure indicated in (3.1) gives a tensor network which is locally indistinguishable from the periodic boundary conditions case $h=v=0$, i.e. it has no defects. Only such cases will be considered below. (When both $h\ \text{mod}\ N$ and $v\ \text{mod}\ M$ are nonzero, this introduces a defect into the network.)
@@
Let $\mathcal{R}$ be a non-rotating RG map with scale factor $b$, and let $h,v$ be two nonnegative integers divisible by $b$, exactly one of which is nonzero. We will say that $\mathcal{R}$ is $(h,v)$-twist friendly, if it preserves the twisted partition function up to reducing all dimensions and twists by the same factor $b$. I.e., for $N,M$ divisible by $b$, we have:
@@
The twist friendliness condition is needed for the validity of some TM spectrum properties stated below. All RG maps known to us, and in particular Gilt-TNR, are $(h,0)$- and $(0,v)$-twist friendly for any $h,v$ divisible by $b$. Intuitively, twist friendliness is a consequence of “locality” of the RG map, and it appears to have been tacitly assumed in TEFR . Rather than trying to formalize the connection to locality, we prefer to state this condition explicitly.
@@ 4.1 LDO definition
Some remarks are in order here. First, note that we posit the existence of an operator $\mathcal{D}$ which acts from the networks with $2\times 2$ holes to networks with $2\times 2$ holes. It’s not obvious that the hole shape may be so preserved. Instead, it can happen that the hole shrinks or expands in size. The shrinking hole case will be discussed in Appendix C, where we will explain some aspects of the results of Lyu:2021qlw ; PhysRevE.109.034111 (see footnote 6). Here in the main text we focus on the $2\times 2$ hole preserving case since this turns out to be relevant for Gilt-TNR.